OpenStudy (foreverandalways13):

whats an equation of a line that passes through (2,-3)? i already have y=-1/2x+4... HELP!!!!!

OpenStudy (anonymous):

y=2x-7

OpenStudy (foreverandalways13):

thanks!!!

OpenStudy (anonymous):

yw

OpenStudy (foreverandalways13):

now put that in point slope form...

OpenStudy (foreverandalways13):

:(....sorry:)

OpenStudy (anonymous):

it is in point slope form, y=mx+b is point slpoe form so it is already in it.

OpenStudy (jdoe0001):

what does y=-1/2x+4 refer to?

OpenStudy (foreverandalways13):

its a line...

OpenStudy (jdoe0001):

well, so, what's the relevance of that line, to "another" crossing on (2, -3)?

OpenStudy (foreverandalways13):

i need it in y-y1=m(x-x1)

OpenStudy (foreverandalways13):

form

OpenStudy (anonymous):

oh hold on

OpenStudy (foreverandalways13):

k

OpenStudy (jdoe0001):

so, the slope-intercept form y=-1/2x+4, and this lines passes through (2, -3)?

OpenStudy (jdoe0001):

this line rather

OpenStudy (foreverandalways13):

yes...i think so...all this gives me is the point:(

OpenStudy (anonymous):

y+3=-1/2(x+2)

OpenStudy (foreverandalways13):

THANKS!!!

OpenStudy (anonymous):

no i mean y+3=-1/2(x-2)

OpenStudy (foreverandalways13):

....again:)

OpenStudy (anonymous):

ur welcome!